in 2 years time susan will be twice the age since she was 5yrs ago how old is she now?
Answer by mananth(16946)

let susan's age be x now
2 years later she will be x+2 years.
..
5 years ago she was x-5 years
..
x+2 = 2(x-5)
x+2=2x-10
2x-x=10+2
x=12 he present age